# ZePT: Zero-Shot Pan-Tumor Segmentation via Query-Disentangling and Self-Prompting

### 07 Dec 2023 v1

@@ Abstract
The long-tailed distribution problem in medical image analysis reflects a high prevalence of common conditions and a low prevalence of rare ones, which poses a significant challenge in developing a unified model capable of identifying rare or novel tumor categories not encountered during training. In this paper, we propose a new Zero-shot Pan-Tumor segmentation framework (ZePT) based on query-disentangling and self-prompting to segment unseen tumor categories beyond the training set. ZePT disentangles the object queries into two subsets and trains them in two stages. Initially, it learns a set of fundamental queries for organ segmentation through an object-aware feature grouping strategy, which gathers organ-level visual features. Subsequently, it refines the other set of advanced queries that focus on the auto-generated visual prompts for unseen tumor segmentation. Moreover, we introduce query-knowledge alignment at the feature level to enhance each query’s discriminative representation and generalizability. Extensive experiments on various tumor segmentation tasks demonstrate the performance superiority of ZePT, which surpasses the previous counterparts and evidence the promising ability for zero-shot tumor segmentation in real-world settings. Codes will be made publicly available.
11footnotetext: Equal contribution. This work was done during Zhongzhen Huang’s internship at Shanghai AI Laboratory.
@@ 1 Introduction
Although OVSS methods have achieved success in segmenting novel categories, their performance heavily depends on the quality of the generated proposals. Our pilot experiments indicate a primary challenge in applying the conventional OVSS strategy to tumor segmentation tasks on medical images: the vision clues of the semantics of tumors are usually more subtle and ambiguous than most common-life objects in natural images, which, as a consequence, makes it difficult to generate high-quality proposals for the unseen tumor categories. Therefore, the commonly adopted assumption in most OVSS methods [12, 34, 16, 65, 46] that the generated proposals cover almost all the potential object-of-interest no longer holds in the scenario of tumor segmentation, where a considerable number of tumors of unseen categories may be poorly covered.
Driven by the aforementioned limitations, we present a novel framework named ZePT for zero-shot tumor segmentation. To generate high-quality proposals that capture unseen tumors, ZePT adopts a query-disentangling scheme that partitions the object queries into two distinct subsets: fundamental queries and advanced queries. Drawing inspiration from the real-world diagnostic procedures of clinical experts, we decouple the learning process into two stages. In Stage-I, we pretrain the fundamental queries via an object-aware feature grouping method to acquire organ-level semantics for precise organ segmentation. In Stage-II, we train advanced queries, guided by self-generated visual prompts emerging from the fundamental queries, to meticulously concentrate on the subtle visual cues associated with tumors. Through its query-disentangling scheme and self-prompting strategy, ZePT explicitly captures fine-grained visual features associated with pathological changes and generates high-quality proposals that precisely cover unseen tumors. At last, we introduce cross-modal alignment between automatically sourced medical domain knowledge and query embeddings to provide weak supervision and augment the model with additional high-level semantic information, further enhancing the model’s generalizability to unseen tumors.
@@ 2 Related Work
Open-Vocabulary Semantic Segmentation. The emerging concept of open-vocabulary semantic segmentation (OVSS) defines a generalized zero-shot semantic segmentation paradigm that allows a model to be trained on conventional vision datasets with close-set labels while possessing the ability to segment an image into arbitrary semantic regions according to text descriptions [12, 34, 16, 65, 46]. Nevertheless, as pointed out in [34], the generated mask proposals trained on base categories are not truly “class-agnostic”. They tend to overfit to seen categories and fail to cover previously unseen obscure objects. This issue hinders the transfer of the power of OVSS to zero-shot tumor segmentation on medical images. Differently, ZePT disentangles the object queries into two sets and adopts a self-prompting strategy to guide the model to explicitly learn semantics related to unseen (novel) tumor categories.
@@
Unseen Lesions/Tumors Detection and Segmentation. Some research efforts [55, 45, 75, 51, 68] also explored innovative approaches to detect or segment unseen lesions/tumors, which formulated the task as out-of-distribution (OOD) detection, making an important step towards safe and robust deep learning methods for clinical practice. Although unseen tumors can be regarded as a kind of OOD sample, the zero-shot segmentation task in this paper is largely different from OOD detection from the perspective that our main objective is to segment and classify multi-class tumors in a zero-shot manner with text descriptions, while OOD detection only recognizes unseen tumors as one single outlier class.
@@ 4 Experiments
Dataset Construction. (1) Training: In Stage-I, we assemble the training sets of $8$ public datasets, including Pancreas-CT [50], AbdomenCT-1K [39], CT-ORG [49], CHAOS [31], AMOS22 [24], BTCV [33], WORD [37] and TotalSegmentator [58]. These datasets exclusively contained organ labels. In Stage-II, we add CT images from the training sets of LiTS [4] and KiTS [19]. The overall seen categories used for training consist of $25$ organ classes and $2$ tumor classes. (2) Inference: We employ the MSD dataset [2] that encompasses a range of segmentation tasks for five tumor types in CTs. Among these, pancreas tumors, lung tumors, colon tumors, and hepatic vessel tumors belong to unseen categories. A real-world, private dataset containing $883$D CT volumes of four distinct colon tumor subtypes is utilized for testing. We follow the data pre-processing in [35] to reduce the domain gap among various datasets. Due to page limits, details of all datasets and pre-processing are described in the supplemental material.
@@ 6.1 Dataset Details
(2) To further evaluate the proposed method, we collect a large real-world colon cancer CT dataset, which consists of 388 patients diagnosed with colon cancer. For each patient, an abdominal CT (venous phase) scan is collected, and the tumor region is annotated by an experienced gastroenterologist and later verified by another senior radiologist. During the annotation phase, the physicians are also provided with the corresponding post-surgery pathological report to narrow down the search area for the tumor(s). All the scans share the same in-plane dimension of $512\times 512$, and the dimension along the z-axis ranges from $36$ to $146$, with a median of $91$. The in-plane spacing ranges from $0.60\times 0.60$ to $0.98\times 0.98$ mm, with a median of $0.76\times 0.76$ mm, and the z-axis spacing is from $5.0$ to $7.5$ mm, with a median of $5.0$ mm. There are four tumor subtypes in this dataset. The full name and incidence count for each disease are shown in Tab. 6. It is important to note that signet ring cell adenocarcinoma and adenosquamous carcinoma constitute only an exceedingly small proportion of all colon cancer cases, illustrating the long-tailed distribution characteristic of real-world disease incidence.

### 22 Mar 2024 v2

@@ Abstract
The long-tailed distribution problem in medical image analysis reflects a high prevalence of common conditions and a low prevalence of rare ones, which poses a significant challenge in developing a unified model capable of identifying rare or novel tumor categories not encountered during training. In this paper, we propose a new Zero-shot Pan-Tumor segmentation framework (ZePT) based on query-disentangling and self-prompting to segment unseen tumor categories beyond the training set. ZePT disentangles the object queries into two subsets and trains them in two stages. Initially, it learns a set of fundamental queries for organ segmentation through an object-aware feature grouping strategy, which gathers organ-level visual features. Subsequently, it refines the other set of advanced queries that focus on the auto-generated visual prompts for unseen tumor segmentation. Moreover, we introduce query-knowledge alignment at the feature level to enhance each query’s discriminative representation and generalizability. Extensive experiments on various tumor segmentation tasks demonstrate the performance superiority of ZePT, which surpasses the previous counterparts and evidences the promising ability for zero-shot tumor segmentation in real-world settings.
11footnotetext: Equal contribution. 🖂Corresponding authors. This work was supported by Shanghai Artificial Intelligence Laboratory. Codes are available at https://github.com/Yankai96/ZePT.
@@ 1 Introduction
Although OVSS methods have achieved success in segmenting novel categories, their performance heavily depends on the quality of the generated proposals. Our pilot experiments in Tab. 1 indicate a primary challenge in applying the conventional OVSS strategy to tumor segmentation tasks: the vision clues of the semantics of tumors are usually more subtle and ambiguous than most common-life objects in natural images, which, as a consequence, makes it difficult to generate high-quality proposals for the unseen tumor categories. Therefore, the commonly adopted assumption in most OVSS methods [13, 35, 17, 67, 48] that the generated proposals cover almost all the potential object-of-interest no longer holds in the scenario of tumor segmentation, where a considerable number of tumors of unseen categories may be poorly covered.
Driven by the aforementioned limitations, we present a novel framework named ZePT for zero-shot tumor segmentation. ZePT adopts a query-disentangling scheme that partitions the object queries into two distinct subsets: fundamental queries and advanced queries. Then we decouple the learning process into two stages, allowing the model to first understand comprehensive organ anatomies and then focus on tumor segmentation, analogous to the learning process of human radiologists. In Stage-I, we pretrain the fundamental queries via object-aware feature grouping to acquire organ-level semantics for precise organ segmentation. In Stage-II, we train advanced queries, guided by self-generated visual prompts emerging from the fundamental queries, to concentrate on the subtle visual cues associated with tumors. Through the query-disentangling and self-prompting, ZePT captures fine-grained visual features associated with pathological changes and generates high-quality proposals that precisely cover unseen tumors. At last, we introduce cross-modal alignment between automatically sourced medical domain knowledge and query embeddings to provide weak supervision and augment the model with additional high-level semantic information, further enhancing the model’s generalizability to unseen tumors.
@@ 2 Related Work
Open-Vocabulary Semantic Segmentation. The emerging concept of OVSS defines a generalized zero-shot semantic segmentation paradigm that allows a model to be trained on conventional vision datasets with close-set labels while possessing the ability to segment an image into arbitrary semantic regions according to text descriptions [13, 35, 17, 67, 48]. Nevertheless, as pointed out in [35], the mask proposals in OVSS methods are not truly “class-agnostic”. They tend to overfit to seen categories and fail to cover previously unseen obscure objects. This issue hinders the transfer of the power of OVSS to zero-shot tumor segmentation on medical images. Differently, ZePT disentangles the object queries into two sets and adopts a self-prompting strategy to guide the model to explicitly learn semantics related to unseen (novel) tumor categories.
@@
Unseen Lesion Detection and Segmentation. Some research efforts [57, 47, 77, 53, 70] also explored innovative approaches to detect or segment unseen lesions/tumors, which formulated the task as out-of-distribution (OOD) detection. Although unseen tumors can be regarded as a kind of OOD sample, the zero-shot segmentation task in this paper is largely different from OOD detection from the perspective that our main objective is to segment and classify multi-class tumors in a zero-shot manner with text descriptions, while OOD detection only recognizes unseen tumors as one single outlier class.
@@ 4 Experiments
Dataset Construction. (1) Training: In Stage-I, we assemble the training sets of $8$ public datasets, including Pancreas-CT [52], AbdomenCT-1K [40], CT-ORG [51], CHAOS [32], AMOS22 [25], BTCV [34], WORD [38] and TotalSegmentator [60]. These datasets exclusively contained organ labels. In Stage-II, we add CT images from the training sets of LiTS [5] and KiTS [20]. The overall seen categories used for training consist of $25$ organ classes and $2$ tumor classes. (2) Inference: We employ the MSD dataset [2] that encompasses a range of segmentation tasks for five tumor types in CTs. Among these, pancreas tumors, lung tumors, colon tumors, and hepatic vessel tumors belong to unseen categories. A real-world, private dataset containing $883$D CT volumes of four distinct colon tumor subtypes is also utilized for testing. We follow the data pre-processing in [36] to reduce the domain gap among various datasets. Due to page limits, details of all datasets and pre-processing are described in the supplemental material.
@@ 6.1 Dataset Details
(2) To further evaluate the proposed method, we collect a large real-world colon cancer CT dataset, which consists of 388 patients diagnosed with colon cancer. For each patient, an abdominal CT (venous phase) scan is collected, and the tumor region is annotated by an experienced gastroenterologist and later verified by another senior radiologist. During the annotation phase, the physicians are also provided with the corresponding post-surgery pathological report to narrow down the search area for the tumors. All the scans share the same in-plane dimension of $512\times 512$, and the dimension along the z-axis ranges from $36$ to $146$, with a median of $91$. The in-plane spacing ranges from $0.60\times 0.60$ to $0.98\times 0.98$ mm, with a median of $0.76\times 0.76$ mm, and the z-axis spacing is from $5.0$ to $7.5$ mm, with a median of $5.0$ mm. There are four tumor subtypes in this dataset. The full name and incidence count for each disease are shown in Tab. 6. It is important to note that signet ring cell adenocarcinoma and adenosquamous carcinoma constitute only an exceedingly small proportion of all colon cancer cases, illustrating the long-tailed distribution characteristic of real-world disease incidence.
@@ 6.4 Additional Ablation Experiments
6.4.6 Methodological clarification of Gumbel-Softmax.
The main motivation of using Gumbel-Softmax is to make the $\operatorname{argmax}$ operation in eq.(4) differentiable [58], where $\operatorname{argmax}$ enables the exclusive one-hot hard assignment of each query. This helps queries focus on distinct visual areas without overlap. We conduct experiments to examine the effectiveness of adopting one-hot hard assignment with Gumbel-Softmax. Our findings reveal that, in comparison to the straightforward use of cross-attention, our choice of incorporating one-hot hard assignment alongside Gumbel-Softmax enhances the DSC by $4.95\%$ on the real-world colon tumor dataset. Also, substituting Softmax with Gumbel-Softmax in equation (2) leads to a $2.86\%$ decrease in DSC on the real-world colon tumor dataset. We assume this occurs because the soft assignment by Softmax, prior to the hard assignment of local features, endows the queries with a global receptive field, allowing them to benefit from long-range contexts.
@@ 6.5 Hyper-Parameter Analysis
6.7 Differences Between ZePT and Existing Zero-Shot Medical Image Segmentation Methods
Research on zero-shot segmentation models is scarcely explored within the medical imaging domain, primarily due to the intricacies involved in medical image segmentation tasks. Early attempts in this area include [41, 4]. Ma et al. [41] proposed a zero-shot CNN model which utilizes two adjacent slices, instead of the target slice, as the input data of deep neural network to predict the brain tumor area in the target slice. This method has the potential to reduce the annotation workload, allowing doctors to only label a subset of the slices. Nevertheless, it requires tumor annotations for training and is capable of segmenting targets only when adjacent slices and their labels are supplied during the training process. Bian et al. [4] introduced an annotation-efficient approach based on zero-shot learning for medical image segmentation. This method leverages the information in data of an existing image modality with detailed annotations and transfer the learned semantics to the target segmentation task with a new image modality. Therefore, their approach more closely resembles Domain Adaptation. These existing methods have yet been definitively proven to have the capability to segment multiple tumors in a strictly "zero-shot" manner. To the best of our knowledge, ZePT represents the first method capable of achieving zero-shot pan-tumor segmentation.